My car has an internal volume of 2600 liters. If the sun heats my car from a temperature of 200 oC to a temperature of 550 oC, what will the pressure inside my car be? Assume the pressure was initially 760 mm Hg
1
Expert's answer
2020-03-02T06:54:56-0500

Use Gay-Lussac's Law( V= const)

"\frac{P_1}{V_1}=\frac{P_2}{V_2}"


"\frac{760 }{220+273}=\frac{P_2}{550+273}"


"P_2= 1269 mmHg"